# Vector-ICL: In-context Learning with Continuous Vector Representations

### v1

@@ A.1 Detailed Experiment Setup
We use two datasets for the text reconstruction task, Parallel Sentence Talks’s English subset [49] and Quora [48].
@@
We train the projectors on the training set of these two datasets and evaluate their performance on the test tests. We measure the reconstruction performance with the BLEU score [43, 44].
@@
We use five datasets for the text classification task. For binary classification, we include IMDB [29], Rotten Tomatoes [42], and the Stanford Sentiment Treebank (SST2) [46]. For multi-class classification, we use the Emotion dataset [45] and the Financial Phrasebank [31]. The binary classification datasets (IMDB, Rotten Tomatoes, and SST2) involve classifying movie reviews as positive or negative. The Emotion dataset classifies Twitter tweets into six categories: anger, fear, joy, love, sadness, and surprise. The Financial Phrasebank categorizes financial news into positive, negative, or neutral sentiments.
@@
We use two datasets for the summarization task, XSum [38] and the English subset of XLSum [17]. They contain newspaper articles and their summaries.
@@
We construct the classification questions around the properties of the underlying text, for example, “Does the sentence contain a proper noun?”, “”Does the input mention anything related to arguing?”. The ground truth is obtained via GPT4o [41] as binary labels. We use the following prompt in ICL in our classification experiments, using one of the example questions:
@@
We use the Chronos [4] time-series Transformers as the encoder. Chronos was pretrained on large scale time-series and is designed to generate the next segments of the time-series. It has proven effective on a wide range of time-series forecasting benchmarks. We take the output of the last time step from Chronos-base as the time-series representation.
We use two datasets for time-series classification, FordA, and FordB, they are both part of the UCR Time Series Classification Archive [12] ranging from 4000 to 5000 time-series for each dataset. We use the following prompt in ICL:
@@
We use Graphormer [59] as the encoder model, specifically the one that was pretrained on large scale quantum chemistry graph datasets [19]. Since the down-stream task (ogbg-molhiv [18]) is a molecule property prediction dataset, and with strong class imbalance (3% positive classes), we finetune the encoder on the training set to provide meaningful baselines and embeddings. We take the output prior to the classification layer of the Graphormer as the graph embedding.
We use the ogbg-molhiv [18] dataset for the graph classification task. ogbg-molhiv is a molecule property prediction dataset consisting of a total 41.12K graphs with node features and edge attributes. It has a strong class imbalance of having around 3% positive class and 97% negative class.
@@
A.2 Text classification and text summarization config in tables
Llama3.1-8B is used as the common LLM and NV-Embed-v1 is used as the text encoder for Vector-ICL. The ICL methods are prompted with up to 50 shots while the soft prompt is trained over the entire training set.
A.3 Additional Experimental Results
@@
A.4 Brain fmri question categories

### v2

@@ Conclusion
Acknowledgement
Our work is sponsored in part by NSF CAREER Award 2239440, NSF Proto-OKN Award 2333790, as well as generous gifts from Google, Adobe, and Teradata. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and should not be interpreted as necessarily representing the views, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for government purposes not withstanding any copyright annotation hereon.
@@ A.1 Detailed Experiment Setup
We use two datasets for the text reconstruction task, Parallel Sentence Talks’s English subset (Tiedemann, 2012) and Quora (Thakur et al., 2021).
@@
We train the projectors on the training set of these two datasets and evaluate their performance on the test tests. We measure the reconstruction performance with the BLEU score (Papineni et al., 2002; Post, 2018).
@@
We use five datasets for the text classification task. For binary classification, we include IMDB (Maas et al., 2011), Rotten Tomatoes (Pang & Lee, 2005), and the Stanford Sentiment Treebank (SST2) (Socher et al., 2013). For multi-class classification, we use the Emotion dataset (Saravia et al., 2018) and the Financial Phrasebank (Malo et al., 2014). The binary classification datasets (IMDB, Rotten Tomatoes, and SST2) involve classifying movie reviews as positive or negative. The Emotion dataset classifies Twitter tweets into six categories: anger, fear, joy, love, sadness, and surprise. The Financial Phrasebank categorizes financial news into positive, negative, or neutral sentiments.
@@
We use two datasets for the summarization task, XSum (Narayan et al., 2018) and the English subset of XLSum (Hasan et al., 2021). They contain newspaper articles and their summaries.
@@
We construct the classification questions around the properties of the underlying text, for example, “Does the sentence contain a proper noun?”, “”Does the input mention anything related to arguing?”. The ground truth is obtained via GPT4o (OpenAI, 2023) as binary labels. We use the following prompt in ICL in our classification experiments, using one of the example questions:
@@
We use the Chronos (Ansari et al., 2024) time-series Transformers as the encoder. Chronos was pretrained on large scale time-series and is designed to generate the next segments of the time-series. It has proven effective on a wide range of time-series forecasting benchmarks. We take the output of the last time step from Chronos-base as the time-series representation.
We use two datasets for the time-series classification task, FordA, and FordB, they are also part of the UCR Time Series Classification Archive (Dau et al., 2019) ranging from 4000 to 5000 time-series for each dataset. We use the following prompt in ICL:
@@
We construct a synthetic pretraining corpus through comprehensive analysis of additional time-series datasets from the UCR Time Series Classification Archive (Dau et al., 2019), specifically MoteStrain, TwoLeadECG, Wafer, PhalangesOutlinesCorrect, and Yoga. Our data curation framework includes multiple statistical approaches: trend identification via linear regression, anomaly detection using z-score analysis, and assessment of temporal stability through variability thresholds. To enhance corpus diversity, we incorporate binary (yes/no) questions targeting specific time-series characteristics. The resulting QA pairs create a rich textual representation that captures the underlying temporal and statistical properties of the data.
@@
We use Graphormer (Ying et al., 2021) as the encoder model, specifically the one that was pretrained on large scale quantum chemistry graph datasets (Hu et al., 2021). Since the down-stream task (ogbg-molhiv (Hu et al., 2020)) is a molecule property prediction dataset, and with strong class imbalance (3% positive classes), we finetune the encoder on the training set to provide meaningful baselines and embeddings. We take the output prior to the classification layer of the Graphormer as the graph embedding.
We use the ogbg-molhiv (Hu et al., 2020) dataset for the graph classification task. ogbg-molhiv is a molecule property prediction dataset consisting of a total 41.12K graphs with node features and edge attributes. It has a strong class imbalance of having around 3% positive class and 97% negative class.
@@
Figure 5: Analyzing LLM’s understanding of projected brain fMRI embeddings after only unsupervised pretraining. We categorize the underlying questions related to the text and measure the mean accuracy for each category, highlighting the LLM’s ability to interpret the embeddings, with only the next token prediction pretraining.
A.2 Case study: what information was perceived from projected brain fmri?
Fig. 5 illustrates the mean accuracy achieved in decoding different categories of brain activity based on fMRI data. The categories, ranging from ”Physical Actions and Movements” to ”Conflict, Urgency, and Change,” represent diverse cognitive and perceptual domains. The grouping and corresponding questions are listed in Table 4.
A.3 Text classification and text summarization config in tables
Llama3.1-8B is used as the common LLM and NV-Embed-v1 is used as the text encoder for Vector-ICL. The ICL and soft-prompt methods are supplied with up to 50 shots as context. The soft-prompt tokens are trained over the entire training set.
A.4 Additional Experimental Results
@@
A.5 Brain fmri question categories